# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

### v1

@@ 2.2. A genericity condition
Suppose that $f(X,Y)\in k[X,Y]$ is a monic polynomial of total degree $d$ over an arbitrary field $k$, such that $f(X,0)$ is squarefree. Cafure and Matera [CM06, §3.1] give an algorithm for testing whether such $f$ has a factor of degree at most $D<d$, which can be improved slightly in this case by setting $\ell_{\max}=dD$ (instead of $2dD$). The correctness of this modified algorithm is verified by the same argument as in the proof of [CM06, Lemma 3.1], similarly to Proposition 2.4 above. Following their notation, we have since the total degrees of $f$ and $g$ are $d$ and $m\leq D$, respectively, the resultant $\mathrm{Res}_{X}(f,g)\in\overline{k}[Y]$ has degree at most $dm$ in $Y$.
Returning to our absolutely irreducible polynomial $f\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, following [CM06, §3.2] and [Kal95, Lemmas 4, 5], there exists a polynomial $\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})$ of total degree at most $2d^{2}$ such that
$\chi(X,Y,Z_{2},\ldots,Z_{n})=f(X+v_{1},w_{2}X+Z_{2}Y+v_{2},\ldots,w_{n}X+Z_{n}Y+v_{n})$
is irreducible in $\overline{\mathbb{F}_{q}}[X,Y,Z_{2},\ldots,Z_{n}]$ with $\chi(X,0,Z_{2},\ldots,Z_{n})\in\overline{\mathbb{F}_{q}}[X]$ squarefree, whenever
$\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})\neq 0.$
The upshot is that (after possibly dividing by the leading coefficient in $X$) we can think of $\chi\in\overline{\mathbb{F}_{q}}(Z_{2},\ldots,Z_{n})[X,Y]$ and apply the aforementioned factorization algorithm. As in [CM06, Theorem 3.3], this produces a polynomial $\Psi_{D}^{(i)}\in\overline{\mathbb{F}_{q}}[Z_{2},\ldots,Z_{n}]$ such that for each root $\alpha_{i,0}\in\overline{\mathbb{F}_{q}}$ of $\chi(X,0,Z_{2},\ldots,Z_{n})$, if $\Psi_{D}^{(i)}(z_{2},\ldots,z_{n})\neq 0$, then $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)=\chi(X,Y,z_{2},\ldots,z_{n})$ has no factors of degree at most $D$ vanishing on $(\alpha_{i,0},0)$. Taking $\Psi_{D}=\prod\Psi_{D}^{(i)}$, we have that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has no factors of degree at most $D$ over $\overline{\mathbb{F}_{q}}$ whenever
$\Xi_{D}(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})=\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})\Psi_{D}(z_{2},\ldots,z_{n})\neq 0.$
To estimate the degree of $\Psi_{D}$, and hence $\Xi_{D}$, the arguments of [CM06, Theorem 3.3] (see also [Kal95, Theorem 5]) remain unchanged, save for the updated value of $\ell_{\max}=dD$. We have
$\deg_{Z_{2},\ldots,Z_{n}}\Psi_{D}^{(i)}$ $\leq\sum_{j=0}^{(D+1)(D+2)/2-1}\ell_{\max}-j$
$=\frac{1}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D\right)$
Multiplying this bound by $d$ yields an upper bound for $\deg_{Z_{2},\ldots,Z_{n}}\Psi_{D}$, and adding $2d^{2}$ to account for $\deg\Upsilon$ gives
$\deg\Xi_{D}\leq\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right).$
Combining this with the well known fact (see e.g. [CM06, Lemma 2.1])
$\{(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}:\Xi_{D}(\mathbf{v},\mathbf{w},\mathbf{z})=0\}\leq(\deg\Xi_{D})q^{3n-3}$
gives the stated result (ii).
The result (i) follows from recognizing that we need only that $\Xi=\Upsilon\Psi_{d-1}^{(1)}\neq 0$ to ensure that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is absolutely irreducible, since at least one factor of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ must vanish on $(\alpha_{1,0},0)$. The corresponding degree bound yields the stated result. ∎
3. Facts about degree 7 forms over $\mathbb{F}_{q}$
Throughout this section, we take $f(x_{0},\ldots,x_{m})$ to be a nonzero homogeneous form of degree $d$ over a finite field $\mathbb{F}_{q}$ and denote by $X_{f}\subset\mathbb{P}^{m}$ the degree $d$ hypersurface cut out by $f=0$. We briefly recall some generalities for degree $d$ forms, before specializing to degree $d=7$.
@@ 3. Facts about degree 7 forms over $\mathbb{F}_{q}$
Suppose $q\geq{d}$ and $f$ is a nonzero degree $d$ form in $m+1\geq 2$ variables. Then $X_{f}(\mathbb{F}_{q})\subsetneq\mathbb{P}^{m}(\mathbb{F}_{q})$.
@@
Suppose $m=1$, i.e. $f$ is a binary form of degree $d$ that vanishes on all $q+1$ points of $\mathbb{P}^{1}(\mathbb{F}_{q})$. Then $x_{0}^{q}x_{1}-x_{0}x_{1}^{q}\mid f$, so we have $\deg f\geq q+1>d$, a contradiction.
We proceed by induction on $m$. If $f$ vanishes on all points in $\mathbb{P}^{m}(\mathbb{F}_{q})$, then $f(x_{0},\ldots,x_{m-1},0)\in\mathbb{F}_{q}[x_{0},\ldots,x_{m-1}]$ is a (possibly now zero) degree $d$ form in $m$ variables that vanishes on all $\mathbb{F}_{q}$-points of $\mathbb{P}^{m-1}$. By the induction hypothesis, it must indeed be the zero form, i.e. $x_{m}\mid f$.
@@
A consequence of Lemma 3.1 is that when $q\geq{d}$ and $H\subset\mathbb{P}^{m}$ is a hyperplane defined over $\mathbb{F}_{q}$, $X_{f}$ vanishes on all $\mathbb{F}_{q}$-points of $H$ if and only if $H\subseteq X_{f}$, i.e. $f$ has a linear factor which corresponds to $H$. Moving forward, we will often conflate these notions.
Let us set a bit more notation. We refer to subvarieties $U\subset\mathbb{P}^{m}$ cut out by the vanishing of a collection of linear forms as linear subvarieties. Given such $U$ defined over $\mathbb{F}_{q}$ of dimension $s$, we use $f|_{U}$ to refer to the form cutting out $X_{f}\cap U$. If $U\subset X_{f}$, then $f|_{U}=0$; otherwise $f|_{U}$ is a nonzero degree $d$ form in $s+1$ variables. Note that $f|_{U}$ is only well defined up to choice of coordinates on $U$ and scaling, but this will not be of concern.
Given linear subvarieties $U,V\subseteq\mathbb{P}^{m}$, their span, denoted $\langle U,V\rangle$, is the minimal linear subvariety of $\mathbb{P}^{m}$ containing $U$ and $V$. In particular, when $U$ and $V$ are $\mathbb{F}_{q}$-points, $\langle U,V\rangle$ denotes the line between them.
For the remainder of the section, $f=f(x_{0},\ldots,x_{m})$ denotes a nonzero degree 7 form in $m+1$ variables over $\mathbb{F}_{q}$. We assume $m>1$.
Lemma 3.2.
Let $q\geq{7}$. Suppose $U,V\subset X_{f}$ are distinct linear subvarieties of dimension $s\geq 0$, defined over $\mathbb{F}_{q}$, such that $\dim(U\cap V)=s-1$. [^1] Throughout, we take the convention that an empty intersection has dimension -1 and codimension $s+1$ in a linear variety of dimension $s$. Then one of the following holds.
  1. (i)
    $f|_{\langle U,V\rangle}$ has an absolutely irreducible factor of multiplicity 1 over $\mathbb{F}_{q}$ of degree at most 3.
  2. (ii)
    $(X_{f}\cap\langle U,V\rangle)(\mathbb{F}_{q})=U\cup V\cup W$ for a linear subvariety $W\subset X_{f}$ (possibly empty) of dimension at most $s$ and defined over $\mathbb{F}_{q}$.
  3. (iii)
    $f|_{\langle U,V\rangle}=0$, i.e. $\langle U,V\rangle\subset X_{f}$.
Proof.
Assume $f|_{\langle U,V\rangle}\neq 0$, i.e. (iii) does not hold. Then $f|_{\langle U,V\rangle}$ is a (nonzero) degree $d$ form in $s+2$ variables, coming from $\langle U,V\rangle\simeq\mathbb{P}^{s+1}$.
$f|_{\langle U,V\rangle}$ has two distinct linear factors $\ell_{U},\ell_{V}$ coming from the fact that it vanishes on $U,V$, since both are codimension 1 in $\langle U,V\rangle$. If either factor has multiplicity 1, then we are in case (i). Otherwise, we write
$f|_{\langle U,V\rangle}=\ell_{U}^{m_{U}}\ell_{V}^{m_{V}}g$
for $m_{U},m_{V}\geq 2$ and $g$ a form of degree $7-m_{U}-m_{V}\leq 3$ over $\mathbb{F}_{q}$.
If $g$ is irreducible over $\overline{\mathbb{F}_{q}}$, we are again in case (i), so assume $g$ is reducible over $\overline{\mathbb{F}_{q}}$. The only way for it to fail to have an irreducible factor of multiplicity 1 over $\mathbb{F}_{q}$ is for $g=\ell_{W}^{m_{W}}$ for $m_{W}=7-m_{U}-m_{V}$, or for $g$ to be irreducible over $\mathbb{F}_{q}$ but factor into conjugate linear forms over $\overline{\mathbb{F}_{q}}$.
In the former case, we have that $f|_{\langle U,V\rangle}$ vanishes on $W\colon\ell_{W}=0$ of dimension $s$. In the latter, we set $W$ to be the (possibly empty) intersection of the linear subspaces over $\overline{\mathbb{F}_{q}}$ cut out by the linear factors of $g$ defined over $\overline{\mathbb{F}_{q}}$. Both of these fall under case (ii). ∎
We would like a version of Lemma 3.2 where (i) implies $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. To achieve this, we need a lower bound for the number of smooth $\mathbb{F}_{q}$-points on a degree $d$ plane curve.
@@
Let q>39q>{39}. Suppose $U,V\subset X_{f}$ are as in Lemma 3.2 with $s\geq 1$. Then one of the following holds.
@@
    $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
@@
    $(X_{f}\cap\langle U,V\rangle)(\mathbb{F}_{q})=U\cup V\cup W$ for $W\subset X_{f}$ linear (possibly empty) of dimension at most $1$ and defined over $\mathbb{F}_{q}$.
  3. (iii)
    $f|_{\langle U,V\rangle}=0$, i.e. $\langle U,V\rangle\subset X_{f}$.
@@
By Lemma 3.2, we have either (ii) or (iii) hold, or that $f|_{\langle U,V\rangle}$ has a simple irreducible factor $g$ of degree at most $3$. In this case, we argue that $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Suppose $\deg g=1$. Then $X_{f}$ has a linear irreducible component $L\simeq\mathbb{P}^{s}$ of multiplicity 1 defined over $\mathbb{F}_{q}$. In particular, $L(\mathbb{F}_{q})=L(\mathbb{F}_{q})^{\mathrm{sm}}=\frac{q^{s+1}-1}{q-1}$. Since $(X_{f}\cap\langle U,V\rangle)-L\subset\mathbb{P}^{s+1}$ is a degree 6 hypersurface, its intersection with $L$ is also a degree 6 hypersurface in $L\simeq\mathbb{P}^{s}$. Such a hypersurface has at most $6\frac{q^{s}-1}{q-1}$ $\mathbb{F}_{q}$-points, so we can find an $\mathbb{F}_{q}$-point in $L(\mathbb{F}_{q})-((X_{f}\cap\langle U,V\rangle)-L)(\mathbb{F}_{q})$, which is smooth in $X_{f}$.
Suppose $f|_{\langle U,V\rangle}$ has no linear components of multiplicity 1. Then the irreducible components of $X_{f}\cap\langle U,V\rangle$ are $U$, $V$, and one cut out by $g=0$, the latter of which is geometrically integral of degree 2 or 3.
Consider first the case of $s=1$, where $f|_{\langle U,V\rangle}$ cuts out a degree 7 plane curve $C$ with a geometrically irreducible component $C^{\prime}$ defined over $\mathbb{F}_{q}$ of degree 2 or 3. Applying Lemma 3.3 with $d=7$, $d^{\prime}=2,3$ and $0\leq g^{\prime}\leq\frac{(d^{\prime}-1)(d^{\prime}-2)}{2}$, we conclude q>19q>{19} suffices to ensure $C_{f}(\mathbb{F}_{q})^{\mathrm{sm}}>0$, hence $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Suppose now that $s>1$. We reduce to the $s=1$ case by an application of Theorem 2.1. After change of coordinates we may assume $U\colon x_{0}=0$ and $V\colon x_{1}=0$ and write
$f=x_{0}^{m_{U}}x_{1}^{m_{V}}g(x_{0},\ldots,x_{m})$
for $m_{U},m_{V}\geq 2$ and $g$ absolutely irreducible of degree 2 or 3. Dehomogenizing by setting $x_{0}=1$ and applying Theorem 2.1(i), when q>39q>{39} there exist $(\mathbf{v},\mathbf{w},\mathbf{z})$ such that $g_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is absolutely irreducible. Now we have
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)=(X-v_{1})^{m_{V}}g_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y).$
Applying the same argument as in the $s=1$ case above (rehomogenizing if necessary), we see that q>39q>{39} is more than sufficient to guarantee the existence of a smooth $\mathbb{F}_{q}$-solution to $f_{\mathbf{v},\mathbf{w},\mathbf{z}}=0$, and hence $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. ∎
We next give an intermediate result in a similar spirit to [LY96, Lemma 5.3], which will be useful later.
@@
Let q>7q>{7}. Suppose $U\subset X_{f}$ is a linear subvariety of positive dimension defined over $\mathbb{F}_{q}$, and $v\in X_{f}(\mathbb{F}_{q})-U$. If there exist distinct codimension 1 linear subvarieties $U_{1}^{(1)},U_{2}^{(1)},U_{3}^{(1)}\subset U$ such that
@@
If $f|_{\langle U,v\rangle}\neq 0$ then $f|_{\langle U,v\rangle}$ has at least four distinct linear factors cutting out the codimension 1 linear subvarieties $U,\langle U_{1}^{(1)},v\rangle,\langle U_{2}^{(1)},v\rangle,\langle U_{3}^{(1)},v\rangle\subset\langle U,v\rangle$. One of these components must have multiplicity 1, so by the argument in the proof of Lemma 3.5 we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. ∎
We conclude this section by examining the situation of a plane curve of degree 7 over $\mathbb{F}_{q}$ which has at least three non-colinear $\mathbb{F}_{q}$-points and no linear components defined over $\mathbb{F}_{q}$. Unfortunately, we cannot guarantee the existence of a smooth $\mathbb{F}_{q}$-point, even when $q$ is sufficiently large. We can, however, find either a smooth $\mathbb{F}_{q}$-point or a line spanned by (singular) $\mathbb{F}_{q}$-points which elsewhere meets the curve at a nonsingular geometric point.
@@
Let q>591q>{591} and suppose $f\in\mathbb{F}_{q}[x_{0},x_{1},x_{2}]$ is a nonzero ternary degree 7 form with no linear factors defined over $\mathbb{F}_{q}$. Suppose further that there exist $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle=\mathbb{P}^{2}$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $u_{1},u_{2}\in X_{f}(\mathbb{F}_{q})$ such that $\langle u_{1},u_{2}\rangle\not\subset X_{f}$ and $X_{f}\cap\langle u_{1},u_{2}\rangle$ contains a geometric point of multiplicity 1.
@@
Suppose $L_{1},L_{2},L_{3}$ meet at a common point $x\in X_{f}(\mathbb{F}_{q})$. Since the points $u,v,w$ are not collinear, we must have that $x$ does not lie on at least one of the lines $\langle u,v\rangle,\langle u,w\rangle,\langle v,w\rangle$. Without loss of generality, assume $x\notin\langle u,v\rangle$. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on the three distinct $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$, so we may take $u_{1}=u$ and $u_{2}=v$.

### v2

@@ 2.2. Genericity conditions
It suffices to prove the statement for $e_{i}=1$ for $1\leq i\leq t$, since $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}=\prod_{1\leq i\leq t}(f_{i})^{e_{i}}_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$.
Let $\Upsilon\in\overline{k}[B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n}]$ be the polynomial obtained from Lemma 2.10. If $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\mathbb{F}_{q}^{2n-1}$ satisfy $\Upsilon(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})\neq 0$, then
$f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})=\prod_{1\leq i\leq t}f_{i}(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$
is squarefree of degree $d$ in $k[X]$.
For $0\leq D_{i}<d_{i}$, and any root $\alpha_{i,0}$ of $f_{i}(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$, we apply Theorem 2.9 to find $\Psi_{i,D_{i}}^{(\alpha_{i,0})}\in\overline{k}[\Delta_{2},\ldots,\Delta_{n}]$ so that whenever $\Psi_{i,D_{i}}^{(\alpha_{i,0})}(\delta_{2},\ldots,\delta_{n})\neq 0$, we have $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no nonconstant factor of degree at most $D$ over $\overline{k}$ vanishing on $(\alpha_{i,0},0)$.
We now address parts (i) and (ii) separately.
  1. (i)
    Set $\Psi=\prod_{1\leq i\leq t}\Psi_{i,d_{i}-1}^{(\alpha_{i,0})}$ which satisfies
    $\deg\Psi=\sum_{1\leq i\leq t}\frac{d_{i}}{8}\left(3d_{i}^{3}-2d_{i}^{2}-3d_{i}+2\right).$
    In this case, if $\Psi(\delta_{2},\ldots,\delta_{n})\neq 0$, we have $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ is irreducible of degree $d_{i}$ over $\overline{k}$ for all $i$, hence $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has type $(d_{1},\ldots,d_{t})$.
  2. (ii)
    Set
    $\Psi=\prod_{1\leq i\leq t}\prod_{\begin{subarray}{c}\alpha_{i,0}\in\overline{k}\\ (f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(\alpha_{i,0},0)=0\end{subarray}}\Psi_{i,D_{i}}^{(\alpha_{i,0})},$
    where in the innermost product, $\alpha_{i,0}$ runs over the $d_{i}$ roots of $f_{i}(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$. This satisfies
    $\deg\Psi\leq\sum_{1\leq i\leq t}\frac{d_{i}D_{i}}{8}\left(-D_{i}^{3}+4d_{i}D_{i}^{2}-6D_{i}^{2}+12d_{i}D_{i}-11D_{i}+8d_{i}-6\right).$
    In this case, if $\Psi(\delta_{2},\ldots,\delta_{n})\neq 0$, we have each $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no nonconstant factor of degree at most $D_{i}$ through any of the $d_{i}$ roots $\alpha_{i,0}$. Thus $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no nonconstant factor of degree at most $D_{i}$.
Note that in either case, while $\Psi$ depends on the choice of $(\boldsymbol{\beta},\boldsymbol{\gamma})$, the upper bound for its degree does not.
In either case, we claim that for at most $(\deg\Upsilon+\deg\Psi)q^{3n-3}$ tuples $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ does $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ fail to satisfy the desired property. Set
$N_{\Upsilon}=\#\left\{(\boldsymbol{\beta},\boldsymbol{\gamma})\in\mathbb{F}_{q}^{2n-1}:\Upsilon(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})=0\right\}$
which is bounded above by $(\deg\Upsilon)q^{2n-2}$ (see e.g. [CM06, Lemma 2.1]). Then the number of $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$ for which $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ fails to have the desired property is bounded above by
$N_{\Upsilon}q^{n-1}+(q^{2n-1}-N_{\Upsilon})\deg\Psi q^{n-2}\leq\left(\deg\Upsilon+\deg\Psi\right)q^{3n-3}.$
Substituting $\deg\Upsilon\leq 2d^{2}$ and the appropriate degree bound for $\Psi$ yields (i) and (ii). ∎
Proof of Theorem 2.2.
When $q>{2d^{2}}$, there exists $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\mathbb{F}_{q}^{2n-1}$ such that $\Upsilon(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})\neq 0$. As in the proof of Theorem 2.1, we define $\Psi\in\overline{k}[\Delta_{2},\ldots,\Delta_{n}]$ appropriately for (i) and (ii) using Theorem 2.9.
If $q>\deg\Psi$, there exists $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ such that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has the same type as $f$ (resp. $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no nonconstant factor of degree at most $D_{i}$). Since we only need the one $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$, the degree of $\Upsilon$ does not contribute to this estimate, and the degree estimates for $\Psi$ alone yield (i) and (ii). ∎
3. Degree $d$ forms over $\mathbb{F}_{q}$
In this section, we record some generalities about degree $d$ forms $f(x_{0},\ldots,x_{n})$ and plane curves defined over $\mathbb{F}_{q}$ before specializing to degrees $d=7$ and $d=11$ in §3.1 and §3.2, respectively.
@@ 3. Degree $d$ forms over $\mathbb{F}_{q}$
Suppose $q\geq{d}$ and $f$ is a nonzero degree $d$ form in $n+1\geq 2$ variables. Then $X_{f}(\mathbb{F}_{q})\subsetneq\mathbb{P}^{n}(\mathbb{F}_{q})$.
@@
Suppose $n=1$, i.e. $f$ is a binary form of degree $d$ that vanishes on all $q+1$ points of $\mathbb{P}^{1}(\mathbb{F}_{q})$. Then $x_{0}^{q}x_{1}-x_{0}x_{1}^{q}\mid f$, so we have $\deg f\geq q+1>d$, a contradiction.
We proceed by induction on $n$. If $f$ vanishes on all points in $\mathbb{P}^{n}(\mathbb{F}_{q})$, then $f(x_{0},\ldots,x_{n-1},0)$ is a (possibly now zero) degree $d$ form in $n$ variables $x_{0},\ldots,x_{n-1}$ that vanishes on all $\mathbb{F}_{q}$-points of $\mathbb{P}^{n-1}$. By the induction hypothesis, it must indeed be the zero form, i.e. $x_{n}\mid f$.
@@
A consequence of Lemma 3.1 is that when $q\geq{d}$ and $H\subset\mathbb{P}^{n}$ is a hyperplane defined over $\mathbb{F}_{q}$, $X_{f}$ vanishes on all $\mathbb{F}_{q}$-points of $H$ if and only if $H\subseteq X_{f}$, i.e. $f$ has a linear factor which corresponds to $H$. Moving forward, we will often conflate these notions.
Let us set a bit more notation. We refer to subvarieties $U\subset\mathbb{P}^{n}$ cut out by the vanishing of a collection of linear forms as linear subvarieties. Given such $U$ defined over $\mathbb{F}_{q}$ of dimension $s$, we use $f|_{U}$ to refer to the form cutting out $X_{f}\cap U$. If $U\subset X_{f}$, then $f|_{U}=0$; otherwise $f|_{U}$ is a nonzero degree $d$ form in $s+1$ variables. Note that $f|_{U}$ is only well defined up to choice of coordinates on $U$ and scaling, but this will not be of concern.
Given linear subvarieties $U,V\subseteq\mathbb{P}^{n}$, their span, denoted $\langle U,V\rangle$, is the minimal linear subvariety of $\mathbb{P}^{n}$ containing $U$ and $V$. In particular, when $U$ and $V$ are $\mathbb{F}_{q}$-points, $\langle U,V\rangle$ denotes the line between them.
Before specializing to $d=7,11$, we record a flexible version of the Hasse–Weil bound for counting smooth $\mathbb{F}_{q}$-points on a plane curve $C$ which allows for $C$ to have singularities or be reducible.
@@
3.1. $d=7$
We now specialize to the case of $d=7$, where $f(x_{0},\ldots,x_{n})$ is a degree 7 form over $\mathbb{F}_{q}$.
@@ 3.1. $d=7$
Let $q\geq{7}$. Suppose $f(x_{0},\ldots,x_{n})$ is a degree 7 form and there exist $k$ distinct linear subvarieties $U_{1},\ldots,U_{k}\subset X_{f}\subset\mathbb{P}^{n}$ of dimension $s$ such that $\dim\langle U_{1},\ldots,U_{k}\rangle=s+1$.
@@
    If $k\geq 4$ then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or $\langle U_{1},\ldots,U_{k}\rangle\subset X_{f}$.
@@
    If $k=3$ and there exists $v\in X_{f}\cap\langle U_{1},U_{2},U_{3}\rangle$ such that $v\notin U_{1},U_{2},U_{3}$, then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or $\langle U_{1},U_{2},U_{3}\rangle\subset X_{f}$.
@@
If $\langle U_{1},\ldots,U_{k}\rangle\not\subset X_{f}$, then $f|_{\langle U_{1},\ldots,U_{k}\rangle}$ is nonzero. In this case, each $U_{i}$ corresponds to a distinct linear factor of $f|_{\langle U_{1},\ldots,U_{k}\rangle}$ defined over $\mathbb{F}_{q}$.
Both statements follow from the fact that when $q\geq{7}$, a degree 7 form with a linear factor of multiplicity one has a smooth solution. If $U_{1}\subset X_{f}\cap\langle U_{1},\ldots,U_{k}\rangle$ corresponds to such a factor of $f|_{\langle U_{1},\ldots,U_{k}\rangle}$, then
$U_{1}(\mathbb{F}_{q})=\frac{q^{s+1}-1}{q-1}\quad\text{and}\quad\#(U_{1}\cap((X_{f}\cap\langle U_{1},\ldots,U_{k}\rangle)-U_{1}))(\mathbb{F}_{q})\leq 6\frac{q^{s}-1}{q-1},$
so there exists $u\in(U_{1}-(X_{f}-U_{1}))(\mathbb{F}_{q})\subset X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}$.
If $X_{f}(\mathbb{F}_{q})=\emptyset$, then $f|_{\langle U_{1},\ldots,U_{k}\rangle}$ has no linear factors of multiplicity one, hence it can have at most three linear factors, proving (i). If it has exactly three such factors, then it must have type $(1^{3},1^{2},1^{2})$. This implies $X_{f}=U_{1}\cup U_{2}\cup U_{3}$, proving (ii) by contrapositive. ∎
We next give an intermediate result in a similar spirit to [LY96, Lemma 5.3].
@@
Let $q\geq{7}$. Suppose $U\subset X_{f}$ is a linear subvariety of positive dimension defined over $\mathbb{F}_{q}$, and $v\in X_{f}(\mathbb{F}_{q})-U$. If there exist distinct codimension 1 linear subvarieties $U_{1}^{(1)},U_{2}^{(1)},U_{3}^{(1)}\subset U$ such that
@@
Let $s=\dim U\geq 1$. If $f|_{\langle U,v\rangle}\neq 0$, we apply Lemma 3.4(i) to $U,\langle U_{1}^{(1)},v\rangle,\langle U_{2}^{(1)},v\rangle,\langle U_{3}^{(1)},v\rangle$ within $\langle U,v\rangle\simeq\mathbb{P}^{s+1}$. If $f|_{\langle U,v\rangle}=0$ then $\langle U,v\rangle\subset X_{f}$. ∎
Next we examine the situation of a plane curve of degree 7 over $\mathbb{F}_{q}$ which has at least three non-colinear $\mathbb{F}_{q}$-points and no linear components defined over $\mathbb{F}_{q}$. Unfortunately, we cannot guarantee the existence of a smooth $\mathbb{F}_{q}$-point, even when $q$ is sufficiently large. We can, however, find either a smooth $\mathbb{F}_{q}$-point or a line spanned by (singular) $\mathbb{F}_{q}$-points which elsewhere meets the curve at a nonsingular geometric point.
@@
Let q>591q>{591}. Suppose $f\in\mathbb{F}_{q}[x_{0},x_{1},x_{2}]$ is a nonzero ternary degree 7 form with no linear factors defined over $\mathbb{F}_{q}$. Suppose further that there exist $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle=\mathbb{P}^{2}$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
@@
Suppose $L_{1},L_{2},L_{3}$ meet at a common point $x\in X_{f}(\mathbb{F}_{q})$. Since the points $u,v,w$ are not collinear, we must have that $x$ does not lie on at least one of the lines $\langle u,v\rangle,\langle u,w\rangle,\langle v,w\rangle$. Without loss of generality, assume $x\notin\langle u,v\rangle$. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on the three distinct $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$, so we may take $y=u$ and $z=v$.